An auto insurance company wants to know the proportion of drivers who always buckle their seat belts when driving. They randomly survey 400 drivers and find that 320 always buckle up. What is the 99% confidence interval for the population proportion of drivers who always buckle their seat belt? Round to 3 decimal places.
Answer #1


Answer Date: 11/11/2019 The 99% confidence interval for the population proportion of drivers who always buckle their seat bel(1-P) 99%C.I. = pzcx = 0.80 +2.576 x 0.80(1-0.80) 400 = 0.748 to 0.852 The 99% confidence interval for the population proport